Two students are 20 years old and one of them decide to leave for space travel

on a rocket made by Tesla, which can travel at a constant speed of 90 percent of

the speed of light. According the time measurement of the traveling student, the

time spent in space was 10 years. Calculate the age of the non-travelling student

when the travelling student travel back to earth.



1
Expert's answer
2021-03-26T08:04:14-0400

"t=t_0+\Delta t=t_0+\frac{\Delta t_0}{\sqrt{1-\frac{v^2}{c^2}}},"

"t=20+\frac{10}{\sqrt{1-\frac{(0.9c)^2}{c^2}}}\approx 43~\text{years}."